OpenStudy (anonymous):

show that the sumation of i(2^i)= 2+n2^(n+2) at i=1 to n+1

OpenStudy (anonymous):

please help! i have chocolate chip cookies

OpenStudy (anonymous):

no

OpenStudy (slaaibak):

\[i * 2^i = 2 +n*2^{n+2}\] Does the Q look like that?

OpenStudy (anonymous):

there is the summation sign in front of the left side

OpenStudy (anonymous):

with i=1 to k+1

OpenStudy (slaaibak):

\[\sum_{i=1}^{n+1} i * 2^i = 2 + n*2^{n+2}\] ?

OpenStudy (anonymous):

yes

OpenStudy (slaaibak):

Cool, ill check it out.

OpenStudy (anonymous):

ok thanks

OpenStudy (slaaibak):

I would use induction. First prove it's true for n=1. Then assume it's true for n. then prove it for n=k+1.

OpenStudy (anonymous):

hy ana

OpenStudy (anonymous):

u there?